W1386 Cold Room is not functional. All liquids were moved to E1386 Cold Room.
The Sorvall RC-6 in W1385 BSTWR is back in service.
BSTWR Elevators are fully functional
The Bio-Rad ChemiDoc was installed last week in E1305B.
E1386B Cold Room cleaning is complete.
E1386B Cold Room will be cleaned by Clean Air Testing Services on Wednesday, December 5, 2012.
E1386B Cold Room will be cleaned on Friday, 12/5/2014.
E1305B Victor II has been repaired.
W1386 dark room film developer is down for repairs due to a mechanical issue.
The W1312A scintillation counter has been shut down. It will be transferred to the Department of Surgery on 9 BSTWR.